Maria has a chronic disease, a kidney disease, and needs dialysis. The cost is $73,000 a year, subject to a $1,000 deductible, 20% co-insurance plus 12 doctor visits at $250 each with a $50 co-pay per visit. If she has a 20% co-insurance and no out-of-pocket maximum, Marie would have to pay $1,000 plus $14,400 plus $600, or $16,000.
This payment is, potentially, $16,000 every year since the disease is chronic. With a $3,000 out-of-pocket maximum, which includes the deductible, co-pays, and co-insurance, the medical expense is capped at how much per year?
1
Expert's answer
2020-03-09T13:36:56-0400


  1. 73 000-1000-72 000 deductible value
  2. 72 000*0.2=14 400 co-insurance
  3. 72 000-14 400+12*250+12*50=72 000-14 400+3000+600=61 200 - total dialysis without deductible, insurance, plus visits to the doctor and surcharge for the visit
  4. 16 000-1 000-600=14 400
  5. 3 000-1 000-600=1400 co-insurance,which will pay in out-of-pocket 
  6. 14 400-1 400=13 000 insurance balance
  7. 16 000-3 000=13 000 insurance balance
  8. 61 200-600=60 600 dialysis balance